If M varies directly as n, and m=9, when n=6, find m when n=15
Answer by stanbon(75887)

If M varies directly as n, and m=9, when n=6, find m when n=15
-------
M = kn
Solve for "k" using m = 9 when n = 6.
------
9 = k*6
k = 3/2
-------
Equation:
M = (3/2)n
---
find m when n = 15
M = (3/2)15
---
M = 45/2 = 22.5
====================
Cheers,
Stan H.
====================